OpenStudy (anonymous):

A 50-L sample of gas undergoes a temperature increase from 40c to 60c. Find new volume.

OpenStudy (rogue):

k = 273 + C 40 C ---> 313 k 60 C ---> 333 k \[\frac {V_1}{T_1} = \frac {V_2}{T_2}\]\[\frac {50 L}{313 k} = \frac { V_2}{333 k}\]\[V_2 = \frac {50 L * 333 k}{313 k} = 53.2 L\]